1. Corleone’s Pizza place makes and delivers pizzas. The cost of ingredients (dough, sauce, cheese, toppings, box) is $4.00 per pie. Corleone pays his employees $50 per day. He also incurs a fixed cost of $80 per day. Recently he has been producing 60 pies per day using 2 workers. Menu price of pizza is $11. (Show work).

  1. Calculate total variable cost per day
  1. Calculate average variable cost per pie
  1. Calculate the average total cost of a pie
  1. Calculate total revenue per day
  1. Calculate total profit per day
  1. Calculate average profit per pie
Answer #1

Total Variable Costs = 60*4+2*50 = 340 per day

Average Variable Costs = Total variable costs/output = 340/60 = 5.67

Fixed Costs = 80

Total Cost = Fixed Cost + Variable Cost

= 340 + 80

TC= 420 per day

Average Total Cost = total cost/output = 420/60 = 7

Total Revenue per day = Price*Quantity = 11*60 = 660

Profits = Total Revenue - Total Costs = 660-420 = 240 per day

Profits per pie = 240/60 = 4